# Bastani, Bastani, and McLaughlin

### v2

@@ 3 Integration to a Policy Optimization Pipeline
Practitioners face two key challenges before they can use Theorem 2.4 to propose a new policy. First, they do not know the true counterfactual outcome distribution and thus must form estimates of the counterfactual means $\hat{\mu}$ and variances $\hat{\sigma}^{2}$ to estimate $\hat{\Pi}_{\Lambda}$. Second, to avoid multiple testing concerns and streamline their decision process, practitioners need to select a single policy from $\hat{\Pi}_{\Lambda}$ to propose and evaluate on $\mathcal{O}$. We suggest practitioners use supervised machine learning algorithms on available training data to form counterfactual estimates and offer a method for selecting $\xi$ to generate $\hat{\pi}^{*}(\xi)$ in terms of the desired minimum $z$-score $Z_{\min}$, and the desired improvement $\lambda$.
@@ 3.1 Estimating the Frontier Using Machine Learning
Forming $\Pi_{\Lambda}$ via Theorem 2.4 requires knowledge about the observational dataset $\mathcal{O}$, namely the observational policy $\pi^{o}$ as well as the mean $\mu$ and variance $\sigma^{2}$ of the counterfactual outcomes. While practitioners often know $\pi^{o}$, the policy which generated $\mathcal{O}$, $\mu$ and $\sigma^{2}$ are fundamentally unknown when optimizing our policy. We suggest first learning estimates of the mean $\hat{\mu}$ and variance $\hat{\sigma}^{2}$ of the counterfactual outcomes in $\mathcal{O}$ using the practitioners choice of machine learning algorithm and an available training dataset $\mathcal{T}$. Our approach can also be applied when $\pi^{o}$ is unknown by additionally estimating $\hat{\pi}^{o}$ from the training dataset.
@@
Once the counterfactual mean $\hat{\mu}$ and variance $\hat{\sigma}^{2}$ estimators are formed, we plug them into Theorem 2.4 with $\mu_{n,t}\leftarrow\hat{\mu}(X^{o}_{n},t)$ and $\sigma^{2}_{n,t}\leftarrow\hat{\sigma}^{2}(X^{o}_{n},t)$ to generate an estimate of $\hat{\Pi}_{\Lambda}$, which contains the Pareto frontier for the counterfactual outcome distribution described by the estimators $\hat{\mu}$ and $\hat{\sigma}^{2}$. It almost certainly underperforms the true frontier of policies related to $\mu$ and $\sigma^{2}$. We engage with this issue more in our technical discussion, but note that as long as $\hat{\mu}$ and $\hat{\sigma}^{2}$ are reasonably close to the true parameters $\mu$ and $\sigma$, $\hat{\Pi}_{\Lambda}$ will extend from the observational policy in a similar direction as $\Pi_{\Lambda}$, ensuring similar policy suggestions for low values of $\xi$.
@@ 3.2 Selecting a Policy from the Optimal Frontier
In addition to generating an estimated frontier, practitioners need to select a policy from the frontier to implement. This problem can be reduced to choosing the value of $\xi$ which generates $\hat{\pi}^{*}(\xi)\in\hat{\Pi}_{\Lambda}$. This can be chosen via guess and check or some form of cross-validation, but this still requires some general idea of $\xi$’s magnitude and may lead to inaccurate evaluation due to concerns with multiple hypothesis testing (Shaffer 1995). Instead, we propose to set $\xi$ using the economic intuition its has as a dual variable.
In the optimization problems defining $\Pi_{\Lambda}$, $\xi$ is the dual variable associated with the constraint $\tau(\pi)=\lambda$. Thus, $\xi$ represents the marginal decrease that the squared standard error $s^{2}(\pi)$ can attain if $\lambda$ is decreased. Thus, to select the optimal policy induced by some $\xi$, we set a limit on how much any policy change can increase the squared standard error in terms of its expected improvement. All expected policy improvements which can be made while introducing less standard error then this limit are included in $\pi^{*}(\xi)$. This intuition offers a conservative way to set $\xi$ for a practitioner hoping to achieve an improvement of $\lambda$ with a $z$-score of at least $Z_{min}$. From the proof of Proposition 2.5, we have that if $\tau(\pi^{*}(\xi))=\lambda$, then $s^{2}(\pi^{*}(\xi))\leq\frac{\xi}{2}\lambda$ due to the limit set by the dual variable $\xi$ and the concavity of $\tau(\pi^{*}(\xi))$ in $\xi$. As a result, we have
$\mathbb{E}[Z(\pi^{*}(\xi))]=\frac{\tau(\pi^{*}(\xi))}{s(\pi^{*}(\xi))}\geq\sqrt{\frac{2\lambda}{\xi}}.$
Thus, we can ensure $\mathbb{E}[Z(\pi^{*}(\xi))]\geq Z_{min}$ by setting
$\xi=\frac{2\lambda}{Z_{min}^{2}}.$
While the limit on $s^{2}(\pi^{*}(\xi))$ is not tight if $\pi^{*}(\xi)\in\Pi_{\text{Opt}}$, we recommend practitioners use this bound as our lack of knowledge with respect to the true counterfactual outcome distribution only allows construction of the inferior $\hat{\pi}^{*}(\xi)$. Using this value of $\xi$ ensures that $\hat{\pi}^{*}(\xi)$ only attempts to make improvements that introduce allowable variance for the given significance level, limiting the concern that noise will drown out significant policy gains.

### v3

@@ 5 Integration to a Policy Optimization Pipeline
Practitioners face two key challenges before they can use Theorem 4.1 to propose a new policy. First, they do not know the true counterfactual outcome distribution and thus must form estimates of the counterfactual means $\hat{\mu}$ and variances $\hat{\sigma}^{2}$ to estimate $\hat{\Pi}_{\Lambda}$. Second, to avoid multiple testing concerns and streamline their decision process, practitioners need to select a single policy from $\hat{\Pi}_{\Lambda}$ to propose and evaluate on $\mathcal{O}$. We suggest practitioners use supervised machine learning algorithms on available training data to form counterfactual estimates and offer a method for selecting $\zeta$ to generate $\hat{\pi}^{*}(\zeta)$ in terms of the desired minimum $z$-score $Z_{\text{min}}$, and the desired improvement $\lambda$.
@@ 5.1 Estimating the Frontier Using Machine Learning
Forming $\Pi_{\Lambda}$ via Theorem 4.1 requires knowledge about the observational dataset $\mathcal{O}$, namely the observational policy $\pi^{o}$ as well as the mean $\mu$ and variance $\sigma^{2}$ of the counterfactual outcomes. While practitioners often know $\pi^{o}$, the policy which generated $\mathcal{O}$, $\mu$ and $\sigma^{2}$ are fundamentally unknown when optimizing our policy. We suggest first learning estimates of the mean $\hat{\mu}$ and variance $\hat{\sigma}^{2}$ of the counterfactual outcomes in $\mathcal{O}$ using the practitioner’s choice of machine learning algorithm and an available training dataset $\mathcal{T}$. Our approach can also be applied when $\pi^{o}$ is unknown by additionally estimating $\hat{\pi}^{o}$ from the training dataset.
@@
Once the counterfactual mean $\hat{\mu}$ and variance $\hat{\sigma}^{2}$ estimators are formed, we plug them into Theorem 4.1 with $\mu_{n,t}\leftarrow\hat{\mu}(X^{o}_{n},t)$ and $\sigma^{2}_{n,t}\leftarrow\hat{\sigma}^{2}(X^{o}_{n},t)$ to generate an estimate of $\hat{\Pi}_{\Lambda}$, which contains the Pareto frontier for the counterfactual outcome distribution described by the estimators $\hat{\mu}$ and $\hat{\sigma}^{2}$. It almost certainly underperforms the true frontier of policies related to $\mu$ and $\sigma^{2}$. We note that as long as $\hat{\mu}$ and $\hat{\sigma}^{2}$ are reasonably close to the true parameters $\mu$ and $\sigma$, $\hat{\Pi}_{\Lambda}$ will extend from the observational policy in a similar direction as $\Pi_{\Lambda}$, ensuring similar policy suggestions for low values of $\zeta$.
The policies trained and evaluated by this pipeline depend on both the training dataset $\mathcal{T}$ and the observed test covariates $X^{o}$. A relevant concern is whether the IPW estimator still provides an unbiased estimate of a policies performance given this information. Under the assumptions of the Rubin Causal model we show that the IPW estimator is still unbiased for policies which depend on a training set and test covariates, $\pi(\mathcal{T},X^{o})$. {repeatproposition}
$\mathbb{E}\left[\hat{\tau}(\pi(\mathcal{T},X^{o}))|\mathcal{T},X^{o}\right]=\tau(\pi(\mathcal{T},X^{o}))$
Thus, while the learned policy itself may vary if the training dataset or test covariates change, IPW still provides an unbiased estimate of the learned policy on the observational dataset.
@@ 5.2 Selecting a Policy from the Optimal Frontier
In addition to generating an estimated frontier, practitioners need to select a policy from the frontier to implement. This problem can be reduced to choosing the value of $\zeta$ which generates $\hat{\pi}^{*}(\zeta)\in\hat{\Pi}_{\Lambda}$. This can be chosen via guess and check or some form of cross-validation, but this still requires some general idea of $\zeta$’s magnitude and may lead to inaccurate evaluation due to concerns with multiple hypothesis testing (Shaffer 1995). Instead, we propose to set $\zeta$ using the economic intuition its has as a dual variable.
In the optimization problems defining $\Pi_{\Lambda}$, $\zeta$ is the dual variable associated with the constraint $\tau(\pi)=\lambda$. Thus, $\zeta$ represents the marginal decrease that the squared standard error $s^{2}(\pi)$ can attain if $\lambda$ is decreased. By selecting $\zeta$, we set a limit on how much any policy change can increase the squared standard error in terms of its expected improvement. All expected policy improvements which can be made while introducing less standard error then this limit are included in $\pi^{*}(\zeta)$. This intuition offers a conservative way to set $\zeta$ for a practitioner hoping to achieve an improvement of $\lambda$ with a $z$-score of at least $Z_{\text{min}}$. From the proof of Proposition 4.2, we have that if $\tau(\pi^{*}(\zeta))=\lambda$, then $s^{2}(\pi^{*}(\zeta))\leq\frac{\zeta}{2}\lambda$ due to the limit set by the dual variable $\zeta$ and the concavity of $\tau(\pi^{*}(\zeta))$ in $\zeta$. As a result, we have
$\mathbb{E}[Z(\pi^{*}(\zeta))]=\frac{\tau(\pi^{*}(\zeta))}{s(\pi^{*}(\zeta))}\geq\sqrt{\frac{2\lambda}{\zeta}}.$
Thus, we can ensure $\mathbb{E}[Z(\pi^{*}(\zeta))]\geq Z_{\text{min}}$ by setting
$\zeta<\frac{2\lambda}{Z_{\text{min}}^{2}}.$
Lower $\zeta$ will raise the probability of achieving a significant result, but lower the expected improvement. While the limit on $s^{2}(\pi^{*}(\zeta))$ is not tight if $\pi^{*}(\zeta)\in\Pi_{\text{Opt}}$, we recommend practitioners use this bound as our lack of knowledge with respect to the true counterfactual outcome distribution only allows construction of the inferior $\hat{\pi}^{*}(\zeta)$. Using this value of $\zeta$ ensures that $\hat{\pi}^{*}(\zeta)$ only attempts to make improvements that introduce allowable variance for the given significance level, limiting the concern that noise will drown out significant policy gains.